### v3

# Computer Vision Datasets and Models Exhibit Cultural and Linguistic Diversity in Perception

Abstract
Computer vision often treats human perception as homogeneous: an implicit assumption that visual stimuli are perceived similarly by everyone. This assumption is reflected in the way researchers collect datasets and train vision models. By contrast, literature in cross-cultural psychology and linguistics has provided evidence that people from different cultural backgrounds observe vastly different concepts even when viewing the same visual stimuli. In this paper, we study how these differences manifest themselves in vision-language datasets and models, using language as a proxy for culture. By comparing textual descriptions generated across 7 languages for the same images, we find significant differences in the semantic content and linguistic expression. When datasets are multilingual as opposed to monolingual, descriptions have higher semantic coverage on average, where coverage is measured using scene graphs, model embeddings, and linguistic taxonomies. For example, multilingual descriptions have on average 29.9 more objects, 24.5 more relations, and 46.0 more attributes than a set of monolingual captions. When prompted to describe images in different languages, popular models (e.g. LLaVA) inherit this bias and describe different parts of the image. Moreover, finetuning models on captions from one language performs best on corresponding test data from that language, while finetuning on multilingual data performs consistently well across all test data compositions. Our work points towards the need to account for and embrace the diversity of human perception in the computer vision community.
1 Introduction
Computer vision researchers often make the implicit assumption that human perception is fundamentally homogeneous. In other words, they assume that—aside from supposedly small and, ultimately, unimportant variations—visual stimuli are perceived similarly by everyone. This homogeneity assumption is visible throughout the vision research field, from datasets to models to metrics. For instance, nearly no commonly used vision dataset, including CIFAR, ImageNet, MS-COCO, and PASCAL-VOC, contain information about their annotator’s background [34], suggesting that it is presumed that this background does not condition their visual perception [27, 68, 98].111Note that annotator information is included in datasets and models from other domains where the annotator’s background is thought to influence annotations, such as in NLP [87, 51] and medical vision [4, 35, 129]. Moreover, vision models trained on data from one demographic of annotators are taken to be representative of the perceptual modes of others across the world [76, 19]. Additionally, many common evaluation metrics average away the supposedly unimportant cross-annotator differences by measuring the mean model performance across multiple ground truth annotations [83, 6].
The homogeneity assumption stands against decades of research in cross-cultural psychology [36]. Early work in psychology suggests that fundamental aspects of visual perception, such as perception of length [101], geometrical intuition [84, 26], and depth [44] vary across people from different cultural and geographic backgrounds. Recent work has shown that differences in cultural backgrounds affect the way that individuals attend to, understand, and talk about visual content [80, 55, 15]. For instance, when looking at the same image, Americans tend to concentrate on the focal objects and its attributes, whereas Japanese tend to focus on the relationships between objects[74]. Other studies show that Westerners tend to perceptually group objects based on category membership, and Easterners based on resemblance [55].
Motivated by evidence from these psychological studies, we investigate if such differences in perception are detectable in vision datasets and model behaviors. We evaluate differences in perception across languages, using language as a rough proxy to culture — this proxy also being used in a strong history of studies on culture [33, 57]. According to the homogeneity assumption, the distribution of semantic content in the descriptions associated with images should be similar. If they are not similar, this would provide further evidence that perceptual biases are observable in vision datasets. Similarly, if models describe the same image differently when prompted to use different languages, we would have evidence that models have inherited such differences in their behavior.
By designing evaluations to examine image captions from datasets, vision-language models, and commercial API across languages, we show that perceptual differences are prevalent in vision datasets and affect model behaviors. In general, for a fixed set of images, multilingual annotations cover more semantic concepts and more expressive diversity than monolingual ones of the same size. Multilingual descriptions have on average more objects, more relations, and more attributes than a set of monolingual captions. Similarly, multilingual descriptions contain more expressive variety; for example, they contain more diversity in abstraction of concepts. These differences are reflected in model behavior: we find that models finetuned on captions translated from one language perform best when evaluated on its corresponding test set, indicating that models internalize the differences between captions in different languages. Overall, our primary contribution is a series of experiments which provide evidence against the homogeneity assumption by demonstrating the diversity (wide and substantive variation) of image annotations across languages and cultures in computer vision datasets and models.
Given these results, we propose several concrete recommendations for computer vision researchers. First, datasets should strive to include annotator information, such as cultural and geographical background. Secondly, multilingual vision models should be preferred over monolingual vision models, as merely translating captions does not guarantee representation of perceptual patterns in the target language. Thirdly, existing multilingual multimodal datasets [130, 98] may be exploited as rich sources of variation in human perception across the globe. Finally, designing computer vision systems explicitly for users from different viewpoints and cultures can improve both performance and accessibility.
2 Related Work
Reductively, there are two general directions in computer vision for understanding and representing diversity across groups of humans. The first direction aims to develop in diverse visual representations by building vision datasets which include images captured from diverse geographical regions around the world [5, 99, 72, 71]. The second direction creates diverse visual knowledge by annotating images with culture- and region-specific information, such as identifying regional dishes, dresses, and ideas [31, 69, 78, 126]. Our work differs from both of these directions in that we investigate differences in the perception of “objective” concepts (rather than culturally specific concepts) across the same images (as opposed to culturally varying images). Our conclusions contest whether any perceived concept is “objective” at all, resulting in implications for how computer vision datasets should be collected and how models should be trained.
In general, human-centric approaches to computer vision center around considerations of human abilities and limitations in the development of models and applications. For example, methods highlighting saliency attempt to identify which image regions and features people find most important [103, 29, 7, 128, 111]. However, they often adopt the homogeneity assumption, suggesting that all people generally find the same types of image regions to be salient. Meanwhile, user-centric vision modeling adapts the models to user-specific preferences and knowledge [105, 23, 117, 92]. However, because these efforts focus on user-specific variations, they do not study patterns across demographics. Our work lies in between these two: we don’t study individual user-behaviors nor average behaviors across all people; instead, we study the differences between groups of people.
Our work is also related to multilingual multimodal modeling, which expands traditional vision models to open-vocabulary language [89, 104] and further to multiple languages [19, 42, 121]. The shift from class labels to image-text pairs [115, 98, 30] opens up computer vision to greater expression of perceptual subjectivity. Although we study differences in multimodal modeling across languages, our main focus is understanding how perception varies between them.
Our work is inspired primarily by research in the social sciences. Psychologists, anthropologists, and philosophers provide strong evidence and reasons that perception is subjective, with broadly ranging studies including cross-cultural psychology [80, 55], usage-based linguistics [116, 62], organizational principles of perception and cognition [123, 54], and the cognitive realities of one’s perceptual experience [124, 39, 108, 43]. Cognitive linguistics posits direct ties between what we say to the way we think and perceive the world [8, 60, 59]. They further suggest that expressed meaning depends on not only what is said—the semantic content of what we say, but also equally on how we say it—the very manner of expression we choose to say it (e.g., specificity of word choice, tone and mood of expression) [62, 119, 9]. That is, a speaker’s conceptualizations has direct influence over what linguistic features or words they reach for and how put them together when they formulate our thoughts into words [61, 107, 25].
3 Experimental Setup
We investigate the following research question: Are there differences in image captions across different languages? We study this at two levels.
1.
Language-level metrics focus on the captions themselves.
Semantic content - measure “what captions (directly) mean.”
Manner of expression - measure “how captions are expressed,” such as the concreteness or directness of language used.
2.
Model-level metrics focus on captions in relation to models.
Model representations - measure differences in caption embeddings.
Model generative outputs - measure the effect of finetuning with certain captions sets on model performance.
Our analysis covers image captions produced in English, French, German, Russian, Chinese, Japanese, and Korean. Together, these 7 languages encompass a wide range of typologically diverse linguistic families. Moreover, the speakers of these languages originate from a large variety of cultures and experiences.
Across each evaluation, if the content measurements are generally similar across captions from different languages, then there are no significant differences in caption content across annotator backgrounds, as suggested by the homogeneity assumption. On the other hand, if these evaluations differentiate captions from different languages, it is evidence against the homogeneity assumption.
3.1 Metrics
We contribute four metrics across two levels to analyze differences in captions across different languages, which we apply across all our caption datasets.
Language-level: Semantic content. The direct semantic content, or meaning, of a caption is captured in what it says about the objects in the scene and the relations between them. Scene graphs are a common representation for this sort of information [47, 20, 45]. Formally, a scene graph is defined as a list of tuples of the form or . Descriptions in our data sources are parsed into scene graphs using FLAN-T5 [22] model fine-tuned on the FACTUAL-MR dataset. We found that LLM-based parsers were better able to resolve complex semantic relationships which often arose in the descriptions than conventional syntax-based parsers [100]. The following metrics can be used to measure the size of a scene graph: object count (number of unique objects mentioned), relationship count (number of object pairs connected by a relationship), and attribute count (number of unique attributes describing objects mentioned).
The union of two scene graphs represents the concepts shared between them. To avoid double counting, in which different words (e.g. “person” and “man”) referring to the same object in the scene, we canonicalize the concepts using WordNet path similarity [75] and cosine similarity between concept embeddings [94] to merge concepts which have different text but ground to the same concept, erring on the side of merging. Multiple relationships between two entities are merged into a single edge.
For our evaluation, we take a scene graph from a caption in one language and generate two scene graphs: a union with two other scene graphs from the same language (monolingual), and a union with two other scene graphs from distinct languages (multilingual). The homogeneity assumption is supported if the monolingual and multilingual scene graphs are roughly the same size, taken across all samples in the dataset. However, if the multilingual scene graphs are larger on average than monolingual scene graphs, then this suggests that captions from different languages contain substantively different semantic content.
Language-level: Manner of Expression. Lexical measures allow us to examine specific dimensions of how text expresses meaning according to principles and frameworks in linguistics. We consider five measures. Concreteness is computed across noun objects and indicates how much a word refers to a perceptible entity [11]. For instance, “purple” has a high concreteness rating (4.04 on a 5-point scale), whereas “beautiful” has a low one (2.16). Text is analytic when it expresses logical and hierarchical thinking patterns, demonstrates clout when it exhibits social status and confidence, is authentic when it is spontaneous, and varies by tone. These last four measures are provided by the Linguistic Inquiry and Word Counts (LIWC) framework [110], and are computed using word dictionaries and aggregation algorithms informed by linguistic and psychological research [10]. Many previous works have demonstrated the use of these measures in the quantification of emotional and psychological expression in text from a variety of contexts [97, 48, 1, 49].
Let be any of these five lexical measures which maps text to a real number. We define the expressive coverage of the set of text as measured by to be . If the mean expressive coverage across multilingual caption sets is larger than over monolingual caption sets, then this suggests – against the homogeneity assumption – that multilingual captions differ in their expression (as measured by ).
Model-level: Model representations. We consider the semantic coverage of a set of text samples, as measured by a text embedding model [94], to be the maximum pairwise cosine distance between embedded text samples on our datasets. If the maximum distance within a multilingual set of captions is larger on average than the maximum distance within a monolingual set of captions with significance, this indicates that captions from the same languages tend to group together in the embedding space relative to captions from other languages.
Model-level: Model generative outputs. We finetune a pretrained vision-language model, GIT [120] – chosen for its simple architecture and robust performance across benchmarks, on 8 constructed image caption sets. These finetuning datasets use the same set of 1.8k training images from the Crossmodal dataset (“XM”) [113]. The first 7 training sets contain translations from a single language (e.g. 1.8k captions all from French, 1.8k captions all from Chinese). The eighth dataset is constructed with equal proportions of captions across all 7 languages. We inference the finetuned models on a hold-out set of 1.8k images from XM and evaluate performance against each of the the corresponding 8 sets of image caption labels. We use SPICE F1-score [3], which measures conceptual overlap over low-level syntax adherence, as our evaluation metric.
We interpret our results using the following logic. Let represent the score of a model fine-tuned on the data from language on the evaluation data from language (suppose bigger is better). The homogeneity assumption is supported if the scores across all evaluations are approximately equal. On the other hand, if , this is evidence against the homogeneity assumption: models learn language-particular concepts when finetuned using data from a specific language, and therefore perform better when evaluated against captions from that language.
3.2 Image caption sources
We perform these evaluations on captions generated at three stages of the computer vision modeling pipeline: by human annotators in datasets, by models, and by commercial APIs for applications.
Datasets. We examine the Crossmodal (“XM”) dataset [114], which contains image descriptions in 36 languages over 3.6k geographically diverse images.
Models. We analyze how models trained with multilingual data describe images across different languages. Specifically, we study LLaVA [70] to produce multilingual captions. Although LLaVA is trained with English data, it inherits probable limited multilingual capabilities from its large language model (LLM) component, LLaMA [118]. In total, we generate 54k captions for 3.6k Crossmodal images in 5 languages with 3 captions each, excluding Japanese and Korean for significant text quality issues. LLaVA allows us to study the significance of language factors across descriptions because the vision representations are constant across text prompts in different languages. See Appendix A.2 for more details on probing multilingual behavior. To provide an English-only baseline, we also compare against descriptions generated by BLIP2 [66] and GIT [120].
APIs. We use the widely adopted Google Vertex API to generate multilingual captions for images – 75.6k captions for 3.6k Crossmodal images in 7 languages, and 42k captions for 2k Visual Genome images. API outputs reflect established industry standards, having gone through extensive quality checks.
3.3 Translation for Fair Comparison
We translate all captions into English with GPT-4 [81] to ensure fair comparison of caption content across languages by eliminating language-based confounders, such as the language-specific parsing of grammatical and expressive structures (for the semantic content and manner of expression evaluations), training data used to produce embeddings (for the model representations evaluation), tokenization (for the model generative outputs evaluation), and so on. Comparing all captions in the same language allows for focus on deeper differences, such as in semantic content and manner of expression. GPT-4 was chosen because we can specify the manner and procedure for translation in the prompt to a degree not available for may other models; specifically, in our case it was more important for translations to be exact and preserve all factual details than to adhere to other translation values and metrics [14]. Several works have demonstrated GPT-4’s competence and control on multilingual tasks [93, 63, 46], particularly in capturing nuances of expression and meaning specific to a language.
Moreover, carrying out our evaluations of captions in their original languages presents significant theoretical and practical challenges. Theoretically, many measurements of linguistic expression are hand-crafted by linguistics and difficult to reproduce in other languages [110]. Other measurements which do have multilingual implementations are not truly language agnostic, so it is dubious if they can be compared across languages. For instance, even advanced multilingual embeddings have been shown to encode language-specific information, especially typology [67, 82, 17, 86]. Practically, implementations of linguistic parsers across languages vary widely in implementation and performance because their design depends on features of that language [90, 16].
We conduct a human evaluation to verify that the translations preserve salient conceptual information. We recruit 15 native speakers to rate 30 translations across each of the non-English languages. Participants mark the translations with an average score of on a 1-to-5 Likert scale, where 5 is ‘entirely accurate / perfect translation” and 4 is “most of the information is preserved in the translation”. Importantly, participants considered of important information in a visual scene (objects, spatial relations, color, etc.) as faithfully preserved across translation. See Appendix A.1 for more details. We also verify statistically that variation among translated captions is not explained by the original language of the captions in Section 5.
4 Language-level experiments
We consider two evaluations of captions at the language-level, corresponding to the semantic content and manner of expression of caption text.
4.1 Results
| Caption Source | Vertex | LLaVA | XM | |||||||
| Metric | Obj.s | Rel.s | Attr.s | Obj.s | Rel.s | Attr.s | Obj.s | Rel.s | Attr.s | |
| Mono | en | 3.65 | 2.96 | 1.67 | 4.54 | 3.79 | 2.75 | 2.59 | 1.54 | 1.27 |
| fr | 3.51 | 2.83 | 1.67 | 5.05 | 4.21 | 3.47 | 2.92 | 1.76 | 1.66 | |
| de | 3.60 | 2.89 | 1.79 | 5.26 | 4.42 | 3.50 | 3.16 | 1.94 | 1.97 | |
| ru | 3.86 | 3.20 | 1.86 | 4.52 | 3.67 | 2.76 | 3.03 | 1.88 | 1.74 | |
| zh.zh,zh | 3.46 | 2.68 | 1.66 | 4.54 | 3.66 | 3.25 | 2.99 | 1.71 | 2.01 | |
| ja.ja.ja | 3.13 | 2.37 | 1.59 | - | - | - | 3.41 | 1.99 | 2.47 | |
| ko | 3.18 | 2.47 | 1.62 | - | - | - | 2.71 | 1.59 | 1.46 | |
| avg | 3.48 | 2.77 | 1.98 | 4.78 | 3.95 | 3.15 | 2.98 | 1.77 | 1.78 | |
| Multi | en,fr,zh | 4.13 | 3.45 | 2.29 | 6.14 | 4.85 | 4.00 | 3.71 | 2.41 | 2.36 |
| fr,zh,ru | 4.24 | 3.48 | 2.40 | 6.15 | 4.76 | 3.99 | 3.92 | 2.57 | 2.59 | |
| de,fr,ru | 4.13 | 3.38 | 2.33 | 6.25 | 4.97 | 4.07 | 3.93 | 2.57 | 2.55 | |
| avg | 4.17 | 3.40 | 2.33 | 5.93 | 4.54 | 3.86 | 4.35 | 2.94 | 2.97 | |
| Caption Source | Vertex | LLaVA | XM | |||||||
| Metric | Conc. | Auth. | Tone | Conc. | Auth. | Tone | Conc. | Auth. | Tone | |
| Mono | en | 1.64 | 23.21 | 1.85 | 2.17 | 35.97 | 6.79 | 1.80 | 33.21 | 8.62 |
| fr | 1.64 | 22.80 | 1.84 | 2.44 | 38.01 | 9.53 | 2.24 | 40.09 | 12.16 | |
| de | 1.67 | 21.68 | 2.03 | 2.53 | 34.33 | 16.64 | 2.11 | 42.22 | 9.74 | |
| ru | 1.66 | 23.07 | 2.63 | 2.27 | 37.20 | 16.48 | 2.03 | 32.35 | 10.90 | |
| zh | 1.51 | 23.51 | 2.05 | 2.33 | 44.63 | 11.56 | 2.26 | 34.02 | 10.59 | |
| ja | 1.50 | 25.01 | 1.96 | - | - | - | 2.25 | 33.83 | 9.18 | |
| ko | 1.56 | 21.67 | 2.05 | - | - | - | 2.14 | 35.28 | 9.07 | |
| avg | 1.60 | 22.99 | 2.06 | 2.35 | 38.03 | 12.20 | 2.12 | 35.86 | 10.04 | |
| Multi | en,fr,zh | 1.75 | 40.16 | 3.98 | 2.54 | 54.94 | 16.74 | 2.08 | 53.12 | 13.78 |
| fr,zh,ru | 1.74 | 36.94 | 4.10 | 2.54 | 54.64 | 20.79 | 2.13 | 54.41 | 14.69 | |
| de,fr,ru | 1.73 | 31.85 | 4.19 | 2.57 | 54.82 | 19.80 | 2.10 | 53.35 | 15.42 | |
| avg | 1.81 | 38.06 | 3.92 | 2.56 | 53.15 | 16.56 | 2.17 | 53.21 | 15.40 | |
Semantic content evaluation. Indeed, we find that multilingual unioned scene graphs cover more objects, relations, and attributes than monolingual ones (Table 1). For instance, unioning the scene graph from an English caption with the scene graphs from a French and a Chinese caption increase the resulting scene graph by objects, relations, and attributes. Further computing the sizes of intersections between monolingual unioned scene graphs (e.g. ) shows that any two languages share only of objects and of relations in common on average (App. Table 12). Since image captions are short and targeted to focus on salient parts of the image, these observed differences between languages likely indicate different tendencies in what counts as salient across caption sets written by humans who speak different languages. Qualitatively, we observe that many of the objects represented in multilingual scene graphs but not in English scene graphs tend to be background objects and objects at varying levels of object composition (e.g. keyboard and CTRL key). Moreover, multilingual scene graphs tend to include more information about perspective (e.g. “zooming”, “fore/background”), color, size, and other details. See Figures 2 and 5 for more examples.
Manner of expression. We find that across all metrics, expressive coverage is generally higher for multilingual rather than monolingual subsets (Table 2). For example, the range of concreteness in caption sets widens by 8.1. This suggests that not only do captions from different languages differ in the logical meaning of their content, but also in how this content is expressed.
4.2 Follow-up Experiments
To further support the suggested implication that captions from different languages mention different content, we do 3 follow-up investigations.
| Monolingual | Multilingual | |||||||||||
| en | de | fr | ru | zh | ja | ko | en,fr,zh | fr,ru,zh | de,fr,ru | avg | ||
| Vertex | 9.6% | 9.7% | 9.9% | 10.3% | 9.8% | 9.9% | 9.3% | 11.7% | 11.7% | 11.2% | 11.5% | |
| 1.55 | 1.60 | 1.62 | 1.69 | 1.60 | 1.59 | 1.49 | 1.92 | 1.92 | 1.84 | 1.90 | ||
| LLaVA | 9.5% | 10.9% | 11.6% | 9.3% | 9.1% | - | - | 12.0% | 12.1% | 12.5% | 12.0% | |
| 2.86 | 3.30 | 3.55 | 2.83 | 2.75 | - | - | 3.71 | 3.70 | 3.86 | 3.69 | ||
“How does the diversity of captions across languages from one model compare with the diversity of captions across multiple models in the same language?” We compare the scene graph sizes when unioning a scene graph from an English caption with the scene graphs from a) two other captions produced by the same model, in different languages; against b) two other captions produced by different models (GIT[120] and BLIP [65]), in the same language. This is an ambitious comparison because different models, even in the same language, are trained with widely varying datasets and procedures. Interestingly, we find that the same-model across-language scene graphs are only slightly smaller than the same-language across-model scene graphs: objects, relations, and attributes. This supports the significance of caption content differences across languages. Furthermore, we find that caption sets generated by one model (Vertex) in different languages have large proportions of concepts not covered by English GIT and BLIP models: of objects, of relations, and of attributes. Interestingly, Asian languages have more such concepts than European ones, which accords with theoretical predictions from cross-cultural psychology [80, 55]. Find further data in the Appendix B.1.
“Do we still observe variation in caption content across languages if we ensure caption concepts are ‘grounded’?” We compare the sizes of multilingual and monolingual concept sets from scene graphs, where all concepts must be matched with an item in the corresponding list of dense annotations documented in the Visual Genome dataset [56]. The conclusion is as before: multilingual scene graphs capture more Visual Genome concepts than monolingual scene graphs (Table 3). For example, unioning an English scene graph with a French and Chinese scene graph captures more Visual Genome-annotated objects than unioning with two other English scene graphs. Interestingly, we also find some evidence that multilingual scene graphs are able to identify objects which are unidentified even in Visual Genome’s dense annotations (App. Table 15).
“Do the observed differences in dataset and model captions correspond to differences in human perception?” Because large-scale image-text datasets do not conduct much annotator information, it is difficult to make detailed and strong inferences about the psychological causes of the observed results, so more work is needed in this direction. However, as a start, we recruited 10 English speakers from the U.S. and 10 Japanese speakers from Japan to caption 30 Visual Genome images and repeated the semantic content evaluation for human-produced captions. See A.3 for study details. We find, in the same pattern as before with model captions, that unioning English scene graphs with Japanese scene graphs expands the size by objects, relations, and attributes over unioning English scene graphs with other English scene graphs. Moreover, a manual inspection of the captions suggests that the captions roughly echo the predictions from cross-cultural perceptual psychology – Japanese captions tend to mention background objects and information more than English ones (see Figures 1 and 10).
5 Model-level experiments
We consider two evaluations at the model-level, corresponding to model representations and model generative outputs.
| Caption Source | Vertex | LLaVA | XM | |
| Mono | en | .19 | .22 | .38 |
| fr | .18 | .29 | .42 | |
| de | .19 | .28 | .43 | |
| ru | .22 | .29 | .40 | |
| zh | .20 | .36 | .46 | |
| ja | .19 | - | .42 | |
| ko | .37 | - | .43 | |
| avg | .22 | .29 | .42 | |
| Multi | en,fr,zh | .37 | .45 | .54 |
| fr,zh,ru | .33 | .47 | .54 | |
| de,fr,ru | .38 | .43 | .49 | |
| avg | .49 | .47 | .52 | |
| Evaluated on | |||||||||
| Fine-tuned on | en | de | fr | ru | zh | ja | ko | multi | |
| en | .23 | .23 | .22 | .22 | .23 | .23 | .23 | ||
| de | .21 | .22 | .22 | .22 | .21 | .23 | .22 | ||
| fr | .25 | .24 | .23 | .24 | .24 | .25 | .25 | ||
| ru | .23 | .23 | .23 | .23 | .24 | .24 | .24 | ||
| zh | .20 | .20 | .20 | .21 | .22 | .22 | .22 | ||
| ja | .21 | .21 | .22 | .21 | .23 | .25 | .22 | ||
| ko | .22 | .22 | .22 | .22 | .24 | .24 | .24 | ||
| multi | |||||||||
Model representations. Indeed, we find that the mean width of caption sets in model representations is larger for multilingual sets than monolingual sets. (Table 4). For instance, the average coverage in embedding space of three English Vertex captions increases by over half from 0.19 to 0.37 when replacing two of them with French and Chinese Vertex captions. We observe a similar pattern in the human-labeled XM dataset, increasing from 0.38 to 0.54. This means that training with captions from multiple languages as opposed to one language may expose the model to a measurably wider range of the representation space.
Model generative outputs. We find that a model fine-tuned on language performs best on language (Table 6), against the homogeneity assumption. We find the same results across models finetuned on data from LLaVA (App. Table 16), and from the human annotated Crossmodal dataset (App. Table 17), including when evaluating with default loss rather than SPICE F-score. Moreover, finetuning on a multilingual collection of captions yields a model which performs consistently well across all evaluation data compositions. In the case of training on human-generated multilingual captions (from the Crossmodal dataset), this model is the second-best model for 5/7 holdout-set languages out of 8 models, coming only behind a model finetuned on that language itself. Together, these suggest that models trained with multilingual data may represent diverse perceptual modes and learn more about the visual world.
One possible confounder is syntactic artifacts introduced during translation. For instance, text translated from German into English might have a unique syntactic structure which distinguishes it from text originally written in English. If this is the case, then it should be possible to identify translated text from one language versus another. To test this limitation, we embed all translated captions using a BERT-based model [94]. We fit a logistic regression model to predict a sample’s original language from these features, and find near-random chance performance at (random chance is ). This suggests that the translation artifact confounder does not explain the observed results.
6 Discussion
Our results show, against the currently prevailing homogeneity assumption in computer vision, that people and models caption images differently in different languages. We conclude with a list of implications, acknowledging limitations with our study, and most importantly, providing a set of recommendations for how researchers should collect vision-language datasets and train models.
6.1 Implications
Reframing the “curse of multilinguality”. We contribute another framing to the ongoing debate related to the “curse of dimensionality.” On one hand, arguments in favor of multilingual language and vision models are often made from an accessibility perspective of expanding information and technology access to individuals [40]. On the other, empirical evidence suggests that increasing the number of languages a model handles degrades monolingual performance [24]. Multilingualism is, therefore, seen as a trade-off against model performance and/or resource cost. Our work provides support for the possibility of a system benefit from including multilingual data. Using organically/human-produced multilingual data introduces substantive changes in the semantic concepts mentioned for the same images. This possibly suggests an “inversion of the curse of multilinguality”: rather than expanding image captioning to multiple languages for accessibility goals at the expense of performance, incorporating the diverse content distribution may improve model robustness [52, 91, 109].
Translating English content to other languages is insufficient. Many efforts to build vision-language models for a non-English language train on translations from English to [76, 19, 32, 18, 85, 88]. Although these systems are usable in , our investigation suggests that they will still reflect an English perceptual bias rather than an “organically ” content distribution (i.e., that native -speakers would provide). Indeed, other works which collect captions from -speakers and train models on them suggest quantitative and qualitative improvements over training from English-to- translated captions [127, 125]. Our work provides an explanation for these observed improvements.
Differences in captions across languages reflect cultural and historical influence. We see the pattern that European languages (English, German, French, and to some extent Russian) are distributed more closely to each other than towards Asian languages (Chinese, Japanese, Korean) in Table 11. Interestingly, besides the geographical proximity within “Western” and “Eastern” languages, the group similarities fall along shared linguistic typological (family) features. Although more research is needed, the current observations align with studies demonstrating linguistic, cultural, and historical influence how we perceive the world [8, 9, 59] and that those influences have a tangible effect on how the people express through language their perceived world [61, 107, 28].
Preventing a “perceptual monoculture”. Our work’s shift in focus from objects to subjects fits within the broader research agenda of aligning with human values and behavior [106, 112]. Importantly, blindly building AI for “humans” in the abstract can substantively neglect populations which are not represented under that signifier, and possibly harm algorithmic equity and fairness goals [102, 12, 38]. The English-dominated language modeling landscape may not only introduce accessibility issues for non-English speakers but also possibly a “perceptual monoculture” [53]. In these ways, our contribution shares motivation and aims with a diverse array of previous works in how models and datasets encode and internalize particular limiting and possibly harmful knowledge, such as by gender [13], nationality [95], by race [96] and language [122, 2].
Perception is not objective. Our results do not just suggest that image captioning involves subjective processes such as emotion and aesthetic judgement [77], or that language is fluid and can represent the same set of concepts in many different ways [73], as others have proposed before. More strongly, it means that even the “objective concepts” within a scene are first and foremost observed by human subjects with particular perceptual tendencies. This aligns with work in the philosophy of consciousness, perception, and language [37, 50, 43].
6.2 Limitations
Using language as a proxy for culture. Our work follows cross-cultural psychology protocol of using language as an approximate proxy for culture [33, 57]. Unfortunately, vision datasets almost never provide cultural or geographical annotator information, making other measurements impossible. Language may be more noisy than other measurements like eye tracking, which has also been used to demonstrate differences between people from different cultures [21]. However, such fine-grained measurements are not available.
Our chosen 7 languages. While our selection of languages is geographically diverse, they are still not representative of global linguistic diversity.
Mid-sized scale. Our experiments operate at a mid-sized scale (thousands of images), emphasizing breadth in languages over depth in images. Future studies may forego such a wide exploration to investigate more specific phenomena at a larger image scale, such as if models differ in their image understanding when trained on captions from different languages. Previous works have shown promise in this direction by showing how better-quality, denser, and more diverse captions can help with better image understanding [79, 58].
Risk of cultural essentialism. Categorizing differences solely with languages may pose a risk of essentializing or stereotyping cultures, suggesting that all members that speak a language perceive the world similarly.
6.3 Recommendations
Dataset construction. Most general image datasets do not provide background information about annotators. Disregarding the cultural background of annotators can unduly overrepresent a set of perceptual behaviors over others. Datasets should also collect annotator background, such as their country or culture, their primary language, and other psychological factors that might substantively affect an individual’s perception. At the minimum, unique annotator identifiers should be included to allow linking of different annotations to their annotator. This information is generally cheap to collect, but will significantly aid the development of perception-aware models and applications.
Multilingual modeling. Currently, many multilingual vision-language models are built by training models on captions translated from English into the desired language. Whenever possible, multilingual vision-language applications should be built by training models collected from native speakers. Our results suggest that the advantage is not only syntactic (i.e. more “natural-sounding” than translations), but that annotations produced in the desired language describe different semantic content and manners of expression.
Multilinguality as perceptual diversity. Multilingual vision-language datasets should be explored as repositories of perceptually diverse annotations. At a basic level, simply training models on multilingual rather than monolingual vision datasets may expose them to a more diverse, rich set of concepts. At a more complex level, specifically designed structured learning paradigms such as contrastive learning may be able to exploit concept diversity across multilingual annotations to yield richer image representations.
Designing for diverse perception. Vision-language applications, such as image captioning, image search, and image generation, should consider, or even condition results on, users’ perceptual background. New systems or models should report metrics across culturally diverse tests sets, and consider running human evaluations with users from culturally diverse backgrounds.
6.4 Conclusion
Similar to evidence presented in social science studies, we provide empirical evidence for cultural and linguistic differences in perception within computer vision datasets and models. Methodologically, we turn the focus from an objective conception of the “ground truth” towards the human subject, which selectively attends to certain objects and qualities in the nearly-infinite field of raw visual features [77, 73]. Our work emphasizes the need for computer vision to understand the cultural [80] and linguistic [64] dimensions of this human subject.
Appendix A Experimental Details
A.1 Translation into English
We prompted GPT-4 [81] to translate text with: “Return the translation (and only the translation) of the following text from [SRC_LANG] into [TGT_LANG] exactly with all details: [TEXT]”.
We find that this prompt produces translations which especially preserve the conceptual details of the original text.
Although some language-specific meanings will inevitably be lost in any translation between languages, we ensure that our English translations are as faithful as possible to the concepts expressed in the original language by conducting a human evaluation. We recruit 2-3 speakers for each of the six non-English languages (French, German, Russian, Chinese, Japanese, Korean), fluent in both the original image and English. Each subject evaluates 30 pairs of original and translated text. Of these 30 pairs, 10 are Vertex captions on Crossmodal images, 10 are LLaVA captions on Crossmodal images, and 10 are Vertex captions on Visual Genome images. This composition ensures wide coverage across image domains and caption format. Each translation evaluation has two parts. Firstly, subjects annotate the overall translation quality on a 1 to 5 scale, in which 1 is “entirely inaccurate”, 2 is “some of the information is preserved”, 3 is “only the most important information is preserved”, 4 is “most of the information is preserved (the translation is adequate but not perfect)”, and 5 is “entirely accurate”. Secondly, subjects examine 11 general categories of concepts in natural visual scenes, provided by TIFA [41]: objects, animals/humans, attributes, activities, spatial relations, counting, food, materials, shapes, locations and colors. Subjects mark each category either as “Good” (the concept was present in the original text and faithfully represented in the translation), “Missed” (the concept was present in the original text but absent or not faithfully represented in the translation), or “N/A” (the concept was not present in the original text). Table 7 demonstrates that the translations are nearly entirely accurate, especially for European languages, and preserve nearly all of the salient content categories for understanding visual scenes.
| Metric | de | fr | ru | zh | ja | ko | |
| Quality Ratings | Mean | 4.95 | 4.76 | 4.82 | 4.63 | 4.48 | 4.48 |
| Median | 5.00 | 5.00 | 5.00 | 5.00 | 5.00 | 5.00 | |
| 25th Percentile | 5.00 | 5.00 | 5.00 | 4.00 | 4.00 | 4.00 | |
| TIFA Categories | Objects | 1.00 | 0.99 | 1.0 | 0.97 | 0.98 | 0.90 |
| Animals/Humans | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | |
| Attributes | 1.00 | 0.89 | 1.00 | 0.93 | 1.00 | 1.00 | |
| Activities | 1.00 | 1.00 | 1.00 | 0.98 | 0.91 | 0.96 | |
| Spatial Relations | 1.00 | 1.00 | 1.00 | 0.92 | 0.94 | 0.96 | |
| Counting | 1.00 | 1.00 | 1.00 | 0.99 | 1.00 | 0.90 | |
| Food | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | |
| Material | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | |
| Shape | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | 1.00 | |
| Location | 1.00 | 0.98 | 1.00 | 0.98 | 0.96 | 0.89 | |
| Color | 1.00 | 0.96 | 1.00 | 0.97 | 1.00 | 1.00 | |
Annotators were allowed to provide free-text explanations for areas in which the translation was inadequate. We provide a random sampling of comments to provide a holistic idea of the translation weaknesses. Overall, the changes to the translations indicated in the comments do not change the content or expression of the text in a substantive way.
| Comment |
| Should use “above” instead of “on” |
| More appropriate to use ”memories” instead of ”impressions” |
| should be ‘small’ balls (remove ‘round’, add ‘small’) |
| Particle suggests that numbers are written “using” sheet of paper, not “on” it. |
| “On the side” is translated as “next to”. |
| “toile d’ araignée” can be directly translated to “cobweb” |
A.2 Probing Multilingual Capabilities in LLaVA
Models like LLaVA which are trained/fine-tuned with English data but which include multilingual LLM components can retain some of these multilingual capabilities. In order to request LLaVA generate captions in a target language, we change the prompt at all levels to correspond to that language language, displayed in Table 9. This works successfully across each of the non-English languages considered in this work, except for Korean and Japanese, which exhibit significantly worse quality.
| Prompt Type | Language | Prompt |
| Roles | English |
(user, assistant) |
| French |
(utilisateur, assistant) |
| German |
(Benutzer, Assistent) |
| System | English |
A conversation between a user and an LLM-based AI assistant. The assistant gives helpful and honest answers. |
| French |
Une conversation entre un utilisateur et un assistant IA basé sur LLM. L’assistant donne des réponses utiles et honnêtes. |
| German |
Ein Gespräch zwischen einem Benutzer und einem auf LLM basierenden KI-Assistenten. Der Assistent gibt hilfreiche und ehrliche Antworten. |
| User Prompt | English |
What is in this image? Answer in English. |
| French |
Qu’est-ce qu’il y a dans cette image? Répondez en français. |
| German |
Was ist auf diesem Bild? Antwort auf Deutsch. |
A.3 Image Captioning User Study
We recruited 10 English speakers from the US and 10 Japanese speakers from Japan (referenced in follow-up to semantic content evaluation). The instructions given to them are presented in Figures 2(a) and 2(b). A sample of the produced captions is given in Table 10.
English | I. |
Two very small boats on a river |
| II. |
Toy boats in the water |
| III. |
A yellow boat and a red boat that appear to be models. |
[…]
Two boats on the water and a building in the back |
| II. |
close-up of a model of a boat and people on the waterfront |
| III. |
Two boats floating on the river and a model of the town in the distance |
English | I. |
Luggage left unattended at a table. |
| II. |
Luggage lined up next to tables with jackets resting on the tables. |
| III. |
luggage sitting next to tables |
[…]
A man sitting in a lobby with lots of suitcases and bags |
| II. |
A man is sitting in a room, and there are several tables filled with luggage nearby. |
| III. |
Man waiting with a lot of luggage |
English | I. |
Cat laying down in an arm chair. |
| II. |
A Siamese cat laying on its back on a couch next to a pillow. |
| III. |
A cat stretched out and upside down on a chair |
[…]
A cat stretches out on a blue chair and a pillow with an embroidered owl next to it. |
| II. |
A cat is relaxing next to a cushion with a picture of an owl on it. |
| III. |
Cat sitting on his back in an armchair with an owl-patterned cushion |
Appendix B Supplementary Data and Figures
Our results across all evaluations are displayed in Table 11.
B.1 Semantics Evaluations
Figure 4 shows that despite an expected diminishing-returns trajectory, continuously unioning even a well-developed existing scene graph with a new language’s scene graph expands it. This suggests that different languages continue to have new information to add to the existing scene graph of visual knowledge. Table 12 displays the sizes of intersections between monolingual scene graphs as measured by the number of objects and relations, using the formula . It is an alternative way to understand the conceptual overlap of different languages. Table 13 shows that scene graphs constructed from captions from the same model but different languages are only slightly smaller than those constructed from captions from the same language but different models. (Referenced in follow-up .) Table 14 shows the intersection sizes between monolingual and multi-model scene graphs. (Referenced in follow-up .) Table 3 shows the sizes of concept sets under intersection with objects documented in Visual Genome. (Referenced in follow-up .) Table 15 shows some samples in which multilingual scene graphs identify objects in the image which are not mentioned in the Visual Genome annotations. (Referenced in follow-up .) Figure 5 shows several examples of scene graphs generated in different languages for different samples.
B.2 Model outputs evaluations
Tables 16 and 17 repeat the same fine-tuning experiment as outlined in the model generative outputs evaluation, but training on LLaVA and XM captions instead of Vertex captions.
| Level | Evaluation | Data Source | Metric | Monolingual | Multilingual | |||||||||
| en | fr | de | ru | zh | ja | ko | en,fr,zh | fr,zh,ru | de,fr,ru | avg222The ‘avg’ column is computed by randomly sampling across all languages. | ||||
| Language | Logical meaning | Vertex | Objects | 3.65 | 3.51 | 3.60 | 3.86 | 3.46 | 3.13 | 3.18 | 4.13 | 4.24 | 4.13 | 4.17 |
| Relations | 2.96 | 2.83 | 2.89 | 3.20 | 2.68 | 2.37 | 2.47 | 3.45 | 3.48 | 3.38 | 3.40 | |||
| Attributes | 1.67 | 1.67 | 1.79 | 1.86 | 1.66 | 1.59 | 1.62 | 2.29 | 2.40 | 2.33 | 2.33 | |||
| LLaVA | Objects | 4.54 | 5.05 | 5.26 | 4.52 | 4.54 | - | 333Japanese and Korean are excluded from LLaVA results due to extremely poor caption generation in the languages. - | 6.14 | 6.15 | 6.25 | 5.93 | ||
| Relations | 3.79 | 4.21 | 4.42 | 3.67 | 3.66 | - | - | 4.85 | 4.76 | 4.97 | 4.54 | |||
| Attributes | 2.75 | 3.47 | 3.50 | 2.76 | 3.25 | - | - | 4.00 | 3.99 | 4.07 | 3.86 | |||
| XM | Objects | 2.59 | 2.92 | 3.16 | 3.03 | 2.99 | 3.41 | 2.71 | 3.71 | 3.92 | 3.93 | 4.35 | ||
| Relations | 1.54 | 1.76 | 1.94 | 1.88 | 1.71 | 1.99 | 1.59 | 2.41 | 2.57 | 2.57 | 2.94 | |||
| Attributes | 1.27 | 1.66 | 1.97 | 1.74 | 2.01 | 2.47 | 1.46 | 2.36 | 2.59 | 2.55 | 2.97 | |||
| Expression | Vertex | Concreteness | 1.64 | 1.64 | 1.67 | 1.66 | 1.51 | 1.50 | 1.56 | 1.75 | 1.74 | 1.73 | 1.81 | |
| Analytic | 0.85 | 0.43 | 0.62 | 1.08 | 2.3 | 2.6 | 2.17 | 2.02 | 2.05 | 1.0 | 2.3 | |||
| Clout | 4.94 | 5.05 | 5.40 | 7.14 | 6.92 | 6.49 | 5.56 | 10.59 | 11.29 | 9.8 | 11.01 | |||
| Authentic | 23.21 | 22.8 | 21.68 | 23.07 | 23.51 | 25.01 | 21.67 | 40.16 | 36.94 | 31.85 | 38.06 | |||
| Tone | 1.85 | 1.84 | 2.03 | 2.63 | 2.05 | 1.96 | 2.05 | 3.98 | 4.10 | 4.19 | 3.92 | |||
| LLaVA | Concreteness | 2.17 | 2.44 | 2.53 | 2.27 | 2.33 | - | - | 2.54 | 2.54 | 2.57 | 2.56 | ||
| Analytic | 2.85 | 4.64 | 14.72 | 7.53 | 23.05 | - | - | 22.03 | 22.29 | 23.76 | 19.61 | |||
| Clout | 14.19 | 19.96 | 15.81 | 23.15 | 21.3 | - | - | 23.06 | 26.78 | 26.84 | 28.72 | |||
| Authentic | 35.97 | 38.01 | 34.33 | 37.20 | 44.63 | - | - | 54.94 | 54.64 | 54.82 | 53.15 | |||
| Tone | 6.79 | 9.53 | 16.64 | 16.48 | 11.56 | - | - | 16.74 | 20.79 | 19.80 | 16.56 | |||
| XM | Concreteness | 1.80 | 2.24 | 2.11 | 2.03 | 2.26 | 2.25 | 2.14 | 2.08 | 2.13 | 2.10 | 2.17 | ||
| Analytic | 1.88 | 2.31 | 1.62 | 0.76 | 5.02 | 5.28 | 5.13 | 4.12 | 4.37 | 4.47 | 4.49 | |||
| Clout | 9.66 | 14.77 | 15.00 | 14.14 | 13.32 | 13.10 | 12.24 | 18.8 | 20.51 | 19.54 | 19.39 | |||
| Authentic | 33.21 | 40.09 | 42.22 | 32.35 | 34.02 | 33.83 | 35.28 | 53.12 | 54.41 | 53.35 | 53.21 | |||
| Tone | 8.62 | 12.16 | 9.74 | 10.90 | 10.59 | 9.18 | 9.07 | 13.78 | 14.69 | 15.42 | 15.40 | |||
| Model | Embeddings | Vertex | Coverage | 0.19 | 0.18 | 0.19 | 0.22 | 0.20 | 0.19 | 0.37 | 0.37 | 0.33 | 0.38 | 0.49 |
| LLaVA | 0.22 | 0.29 | 0.28 | 0.29 | 0.36 | - | - | 0.45 | 0.47 | 0.43 | 0.47 | |||
| XM | 0.38 | 0.0.42 | 0.43 | 0.40 | 0.46 | 0.42 | 0.43 | 0.54 | 0.54 | 0.49 | 0.52 | |||
| Fine-tuning | Vertex | Best Model444Each cell value is the original language of the caption set used to train the model which performed best as evaluated by the caption set originally from the language specified in the column. | en | de | fr | ru | zh | ja | ko | - | - | - | multi | |
| LLaVA | en | de | fr | ru | zh | - | - | - | - | - | multi | |||
| XM | en | de | fr | ru | zh | ja | ko | - | - | - | multi | |||
| en | de | fr | ru | zh | ja | ko | |
| en | 3.65 / 2.96 | 2.34 / 1.20 | 2.41 / 1.26 | 2.39 / 1.24 | 2.15 / 0.97 | 2.04 / 0.91 | 2.02 / 0.89 |
| de | 3.51 / 2.83 | 2.37 / 1.24 | 2.47 / 1.32 | 2.14 / 0.98 | 2.04 / 0.91 | 2.06 / 0.94 | |
| fr | 3.60 / 2.89 | 2.44 / 1.27 | 2.16 / 0.97 | 2.07 / 0.91 | 2.09 / 0.95 | ||
| ru | 3.86 / 3.2 | 2.25 / 1.06 | 2.13 / 0.98 | 2.12 / 0.96 | |||
| zh | 3.46 / 2.68 | 2.08 / 0.95 | 2.04 / 0.92 | ||||
| ja | 3.13 / 2.37 | 2.10 / 1.02 | |||||
| ko | 3.18 / 2.47 |
| en,fr,zh | fr,de,ru | multi-model | ||
| Vertex | Objects | 4.31 | 4.25 | 4.63 |
| Relations | 3.60 | 3.56 | 3.64 | |
| Attributes | 2.13 | 2.15 | 2.19 | |
| LLaVA | Objects | 5.87 | 6.02 | 6.65 |
| Relations | 4.84 | 4.97 | 5.42 | |
| Attributes | 4.10 | 4.07 | 2.88 |
| Language | ||||||||
| en | de | fr | ru | zh | ja | ko | mm | |
| Objects | 1.96 | 1.92 | 1.93 | 1.97 | 1.85 | 1.73 | 1.76 | 3.59 |
| Relations | 0.79 | 0.76 | 0.74 | 0.78 | 0.70 | 0.62 | 0.64 | 2.51 |
| Attributes | 0.44 | 0.36 | 0.37 | 0.42 | 0.37 | 0.37 | 0.33 | 1.45 |
| Image | VG Objects | Scene Graph Objects |
woman, sign, man, bag, license plate, car, person, leg, satchel | umbrella, sandwich restaurant, street, rain |
woman, key, notes, page, keyboard, pencil case, laptop, student | table |
leaves, sign, sky, cloud, trees, roof, train, steam cloud, ground, lamp, green leaves, cables, pole, tracks, locomotive, train car, tree, steeples, gravel, steam, bush, door, wheel | number, logo, inscription |
tray, writing, cloth, stove door, light, oven back, bird necklace, mitt, shirt, apron, stove, burner, strings, aprontop, towel, board, pizza, shortsleeveshirt, menu, woman, necklace, pizzas, pan, oven, sheet | chalkboard |
giraffe tail, spot, rock, giraffe, rocks, grass |
bird (left of image) |
| Evaluated on | ||||||
| en | de | fr | ru | zh | multi | |
| en | 0.225 | 0.229 | 0.219 | 0.218 | 0.230 | |
| de | 0.213 | 0.219 | 0.217 | 0.215 | 0.219 | |
| fr | 0.248 | 0.240 | 0.234 | 0.236 | 0.246 | |
| ru | 0.226 | 0.234 | 0.228 | 0.231 | 0.239 | |
| zh | 0.199 | 0.202 | 0.199 | 0.207 | 0.216 | |
| multi | ||||||
| Evaluated on | ||||||||
| en | de | fr | ru | zh | ja | ko | multi | |
| en | 0.124 | 0.1421 | 0.120 | 0.114 | 0.129 | 0.130 | 0.148 | |
| de | 0.158 | 0.152 | 0.143 | 0.124 | 0.140 | 0.146 | 0.149 | |
| fr | 0.182 | 0.142 | 0.143 | 0.130 | 0.146 | 0.150 | 0.154 | |
| ru | 0.172 | 0.136 | 0.152 | 0.125 | 0.137 | 0.142 | 0.148 | |
| zh | 0.144 | 0.116 | 0.129 | 0.120 | 0.130 | 0.142 | 0.130 | |
| ja | 0.144 | 0.128 | 0.137 | 0.125 | 0.124 | 0.144 | 0.135 | |
| ko | 0.151 | 0.116 | 0.131 | 0.116 | 0.115 | 0.134 | 0.134 | |
| multi | ||||||||
![[Uncaptioned image]](extracted/5459933/imgs/user-study-example-panel/example_caption_image_user_study2.png)
![[Uncaptioned image]](extracted/5459933/imgs/user-study-example-panel/example_caption_image_user_study3.png)
![[Uncaptioned image]](extracted/5459933/imgs/user-study-example-panel/example_caption_image_user_study4.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example8.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example2.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example3.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example4.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example7.png)